29. A solution that contains 800 mg/dL serum is diluted 1⁄2, then 1⁄4, then 1/8, and then again 1/16 in a series. Determine the following:
a. Final dilution:
b. Final concentration:
30. A hepatitis A antibody titer analysis was executed on a blood sample. The dilutions tested were 1⁄4, 1/8, 1/32, and 1/64. A test for the presence of an antibody was then made by adding a hepatitis A antigen to each of the tubes. Positive reactions were seen in the 1⁄4, 1/8, and 1/32 dilutions and negative in the 1/64 dilution. Find the antibody titer.
1
Expert's answer
2020-11-17T10:45:16-0500

29.

1/2. 800/2 = 400. mg/dL

1/4. 400/4 = 100. mg/dL

1/8. 100/8 = 12.5. mg/dL

1/16. 12.5/16 = 0,78125 mg/dL.

a. Final dilution: 1/1024.

b. Final concentration: 0,78125 mg/dL.

30.

To calculate antibody titer, a blood serum sample containing antibody is diluted in serial ratios (1:2, 1:4, 1:8, 1:16...etc.). Using an appropriate detection method (e.g., colorimetric, chromatographic, etc.), each dilution is tested for the presence of detectable levels of antibody. The assigned titer value is indicative of the last dilution in which the antibody was detected.

In this case antibody titer is 32.